If a sequence is defined recursively by f(1)= -2 and f(n-1) = 3(n-1)-4, then find the first 4 terms of the sequence?
Answer by ikleyn(52786)

.

            The problem formulation, strictly saying, is incomplete, since it DOES NOT define for which values of "n" the given formula is valid.

            But if we assume that it is valid for any integer n >= 3, then see my solution below.

            Read it VERY ATTENTIVELY.

            Also notice that the formula is INVALID for n= 2.


Solution

It is INCORRECT to call this formula recursive.


There is NO recursion in it.


It is an EXPLICIT formula  f(m) = 3m-4, for any integer m >= 2, if we accept my assumption above.


So, f(2) = 3*2 - 4 = 6 - 4 = 2;

    f(3) = 3*3 - 4 = 9 - 4 = 5;

    f(4) = 3*4 - 4 = 12 - 4 = 8.


You can easily calculate million next terms in this way, if you want.

--------------

Very important  POST-NOTE :


        It is a  GROSS  MISTAKE  to call this formula recursive (!)